# MoReMouse: Monocular Reconstruction of Laboratory Mouse

### v1

@@ 3.2.1 Transformer and Triplane Representation
Figure 4: Architecture of the MoReMouse reconstruction network. Given a monocular input image, a DINO-V2 encoder extracts high-level image features, which are processed by a transformer-based decoder to generate a triplane representation. These triplanes, combined with learnable positional embeddings, are queried by a Multi-Head MLP to predict per-point color, density, and geodesic embedding. Rendering is performed via either NeRF (for volumetric supervision) or DMTet (for explicit geometry output), producing RGB images, color-coded embeddings, and surface meshes.
3.2.2 Feature Embedding for Surface Consistency
The reconstruction of mice poses significant challenges due to their black fur, which lacks distinctive surface features, and the presence of intricate fine-scale geometric details. Inspired by Animatable Gaussians [20], we incorporate feature embeddings into the mouse surface representation to enhance reconstruction stability. These embeddings serve as additional color channels during training, acting as supervisory signals for MoReMouse rendering.
Specifically, given the set of mouse model vertices $V\in\mathbb{R}^{13059\times 3}$, we compute the pairwise geodesic distances between vertices to obtain the geodesic distance matrix:
$G_{V}\in\mathbb{R}^{13059\times 13059},\quad G_{V}(i,j)=d_{\text{geo}}(V_{i},V_{j}),$ (6)
@@ 3.2.2 Feature Embedding for Surface Consistency
To learn a continuous correspondence embedding that captures the intrinsic structure of the mouse surface, we define the embedding matrix $E\in\mathbb{R}^{13059\times 3}$ and optimize it by minimizing the $L_{2}$ loss between the embedding-derived distance matrix $G_{E}$ and the geodesic distance matrix $G_{V}$:
@@
Notably, a simple rotation of the embeddings in color space can generate new feature mappings:
$E^{\prime}=R(\theta)E,$ (9)
where $R(\theta)$ is a rotation matrix. Extending the embeddings to higher dimensions could further improve reconstruction robustness. However, we leave the exploration of such extensions to future work.
3.2.3 Training and Optimization
Following the method outlined in Section 3.1.2, we generate freely renderable Mouse Avatars for each frame of the ”markerless_mouse_1” video sequence. As the dataset consists of 18,000 frames, we use the first 6,000 frames for training and the last 6,000 frames for testing. To construct the training set, we sample two sets of 64 viewpoints per frame, rendering 12,000 multi-view scenes in total. The optical axes of these viewpoints converge at the scene origin.
@@ 3.2.3 Training and Optimization
We employ a two-stage training strategy, where NeRF-based volumetric rendering is first trained for 60 epochs to ensure proper density field convergence, followed by DMTet-based explicit geometry modeling for an additional 100 epochs to refine surface details.
The total training loss is defined as:
@@
4 Experiments
4.1 Experimental Setup
Synthetic Evaluation Data.
We use the last 6,000 frames from the markerless_mouse_1 sequence to generate a synthetic multi-view dataset consisting of 4 orthogonal viewpoints. These views are rendered using our Gaussian avatar pipeline and serve as a controlled evaluation set where ground-truth camera parameters and rendering conditions are known.
Real Captured Data.
To assess performance under realistic conditions, we collect a real-world multi-view dataset using industrial cameras. This dataset contains 5,400 frames of freely moving mice captured from four calibrated viewpoints. The camera rotation angles relative to the reference view are $0^{\circ}$, $-33.39^{\circ}$, $-78.60^{\circ}$, and $-181.02^{\circ}$, respectively.
Baselines.
  • TripoSR [39]: A state-of-the-art open-source Large Reconstruction Model (LRM) known for strong single-view 3D generation performance.
  • Triplane-GS [50]: A LRM variant based on Gaussian representations.
  • LGM [38]: A multi-view diffusion-based method that reconstructs 3D Gaussian point clouds.
  • InstantMesh [45]: A mesh-based 3D reconstruction method using multi-view diffusion priors.
For synthetic data, we additionally evaluate the multi-view diffusion baselines (LGM and InstantMesh) using ground-truth views as input, which serves as an approximate upper bound for multi-view reconstruction.
Evaluation Protocol.
All methods are evaluated on the task of novel view synthesis, since obtaining ground-truth 3D geometry for mouse subjects is impractical. For each frame, a single view is used as input, and the remaining three views are used to evaluate the visual fidelity of synthesized images. We report three standard image-based metrics: Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM), and Learned Perceptual Image Patch Similarity (LPIPS).
4.2 Experimental Results
Quantitative Results.
Table 1: Quantitative results on both synthetic and real-world datasets. MoReMouse achieves the best performance across all metrics in both domains, significantly outperforming baseline methods in terms of reconstruction fidelity.
Dataset Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
synthetic Ours 22.027 0.9660 0.05279
TripoSR 13.673 0.8032 0.18255
Triplane-GS 18.049 0.9268 0.10151
LGM 14.460 0.8805 0.13274
InstantMesh 15.821 0.8987 0.11312
real Ours 18.422 0.9478 0.08674
TripoSR 11.518 0.8114 0.19672
Triplane-GS 16.789 0.9298 0.11002
LGM 15.215 0.9197 0.12838
InstantMesh 15.631 0.9175 0.11339
Across all three metrics for novel view synthesis, MoReMouse consistently outperforms the baseline methods, demonstrating its ability to effectively infer plausible mouse geometry and pose from a single image. Notably, in Table 2, MoReMouse uses only a single input view, while LGM and InstantMesh are evaluated with 4 and 6 views respectively. For LGM, the 4 input views happen to coincide with the 4 evaluation viewpoints, offering an idealized setting that may favor the method. Nonetheless, MoReMouse achieves comparable performance despite relying on significantly less input information, underscoring its strong generalization capabilities in monocular reconstruction scenarios.
We also observe that the quantitative metrics on the real-world dataset are generally lower than on the synthetic dataset, especially for PSNR. This is largely due to the camera-mouse configuration: in the real dataset, the cameras are fixed while the mouse moves, whereas MoReMouse and all baselines assume that the mouse is centered and aligned across input-output views. This mismatch introduces pixel-level alignment errors, reducing PSNR. However, the perceptual metrics SSIM and LPIPS remain high, indicating that MoReMouse still reconstructs plausible and correct mouse poses and shapes. This confirms that models trained on synthetic data can generalize to real-world mouse motion capture scenarios.
Table 2: Comparison with multi-view diffusion-based methods using ground-truth viewpoints on the synthetic dataset. Despite using only a single view, MoReMouse performs competitively with multi-view approaches that use 4 or 6 calibrated inputs.
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Ours (single view) 22.027 0.9660 0.05279
LGM (4 views) 23.998 0.9698 0.06458
InstantMesh (6 views) 28.295 0.9831 0.02716
Qualitative Results.
To visually compare MoReMouse with baseline methods, we select three representative frames from both the synthetic and real-world sequences. As shown in Fig. 5 and Fig. 6, MoReMouse generates mice with coherent appearance and physically plausible postures, accurately capturing tail and limb movements. These results remain stable even though each frame is inferred independently, showing smooth motion continuity.
In contrast, baseline methods often exhibit structural distortions in novel views and inconsistent pose transitions across frames. Diffusion-based approaches such as InstantMesh and LGM tend to produce view-inconsistent geometry, likely due to the absence of strong priors on mouse anatomy. Single-view approaches like Triplane-GS and TripoSR frequently fail to reconstruct accurate poses, especially under extreme articulations or occlusions.
4.3 Ablation Study
4.3.1 Comparison with Finetuned TripoSR
In previous experiments, we compared MoReMouse with four general-purpose single-view 3D reconstruction baselines, demonstrating the effectiveness of our complete training pipeline. However, these reconstruction models were pre-trained on datasets that may not contain sufficient small animal or mouse-like samples, which could limit their performance on our mouse-specific benchmarks.
To assess whether the improvements of MoReMouse are due to architectural and representation innovations rather than only data augmentation, we fine-tune TripoSR on our mouse dataset using a similar training protocol. Specifically, we train the model for 120 epochs with a learning rate of $1\times 10^{-6}$. As shown in Table 3 and illustrated in Fig. 7, the fine-tuned TripoSR achieves improved performance compared to its original version, but still lags behind MoReMouse across most metrics. This suggests that our model architecture, including the triplane encoding and geodesic feature embedding, plays a critical role in achieving high-fidelity reconstructions.
Dataset Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
synthetic Ours 22.027 0.9660 0.05279
TripoSR-Tuned 21.996 0.9676 0.06333
real Ours 18.422 0.9478 0.08674
TripoSR-Tuned 18.162 0.9461 0.09354
4.3.2 Effect of Feature Embedding
Figure 6: Qualitative comparison of novel view synthesis results on real data. MoReMouse generalizes well to real-world mouse recordings, maintaining consistent geometry and appearance across views, while baseline methods often suffer from inconsistent shapes and misaligned postures.
Dataset Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
synthetic full model 22.027 0.9660 0.05279
w/o embedding 21.805 0.9655 0.05467
real full model 18.422 0.9478 0.08674
w/o embedding 18.250 0.9469 0.08767
4.4 Failure cases
Figure 9 showcases typical failure cases on real-world data. The first row highlights the model’s inability to generalize to rare upright postures, due to the absence of such poses in the training set. This leads to implausible geometry and emphasizes the need for more pose-diverse avatar data.
Figure 9: Representative failure cases. Top: failure due to lack of standing mouse data in training, leading to unrealistic reconstructions. Bottom: failure due to tail-induced occlusion from shaved fur, which interferes with head feature extraction.
5 Discussion
Summary
In this paper, we present MoReMouse, the first monocular dense 3D reconstruction model tailored for laboratory mice. Unlike existing single-view 3D methods designed for general objects or humans, MoReMouse addresses the unique challenges posed by small animals, including non-rigid deformation, low surface texture, and limited training data. Our approach integrates transformer-based triplane representations, a Gaussian avatar modeling pipeline, and geodesic surface feature embeddings to achieve robust and high-fidelity reconstructions from a single image. We demonstrate strong performance across both synthetic and real datasets, outperforming state-of-the-art reconstruction baselines in terms of visual quality and perceptual consistency. Ablation studies further validate the architectural contributions of our model beyond dataset-specific priors. We hope that MoReMouse can serve as a foundation for future research in computational ethology and non-rigid animal modeling, promoting more precise, efficient, and scalable behavioral analysis tools for biomedical research.
Limitations.
Ethical Statement.
Acknowledgements
Supplementary Material
1 MoReMouse Model Details
We summarize the architectural and training hyperparameters of the MoReMouse model in Table 5, covering key components such as the image encoder, triplane representation modules, multi-head MLP, NeRF and DMTet renderers, and training configurations.
Table 5: MoReMouse Architecture and Training Hyperparameters
@@ 1 MoReMouse Model Details
Hidden layers for density/feature/deformation heads $1$
@@
Density bias $-1.0$
Renderer-DMTet SDF bias $-4.0$
@@
$\lambda_{\text{mse}}$, $\lambda_{\text{LPIPS}}$ $1.0$
@@
$\lambda_{\text{smooth}}$, $\lambda_{\text{depth}}$ $0.2$

### v2

@@ 2 MoReMouse Architecture
Figure 4: Overview of the MoReMouse architecture. Given a single image, a DINOv2 encoder and transformer decoder generate triplane features, which are queried to produce color, density, and embeddings. Rendering is performed via NeRF or DMTet for volumetric or surface outputs.
Feature Embedding for Surface Consistency
Mouse reconstruction is challenging due to black fur with weak texture cues and fine-scale surface details. Inspired by Animatable Gaussians (li2024animatable), we incorporate feature embeddings computed from geodesic distances (shamai2017geodesic) into the surface representation. These embeddings are used as additional color channels during training, serving as semantic priors to enhance reconstruction stability.
As shown in Fig. 2, given the set of mouse model vertices $V\in\mathbb{R}^{N_{V}\times 3}$, we compute the pairwise geodesic distances between vertices to obtain the geodesic distance matrix:
$G_{V}\in\mathbb{R}^{N_{V}\times N_{V}},\quad G_{V}(i,j)=d_{\text{geo}}(V_{i},V_{j}),$ (5)
@@
To learn a continuous correspondence embedding that captures the intrinsic structure of the mouse surface, we define the embedding matrix $E\in\mathbb{R}^{N_{V}\times 3}$ and optimize it by minimizing the $L_{2}$ loss between the embedding-derived distance matrix $G_{E}$ and the geodesic distance matrix $G_{V}$:
@@
Training and Optimization
@@
We employ a two-stage training strategy, where NeRF-based volumetric rendering is first trained for 60 epochs to ensure proper density field convergence, followed by DMTet-based explicit geometry modeling for an additional 100 epochs to refine surface details. Details of the training loss terms and hyperparameters are provided in the supplementary material.
4 Experiments
1 Experimental Setup
Synthetic Evaluation Data.
We use the last 6,000 frames from the “markerless_mouse_1” sequence to generate a synthetic multi-view dataset consisting of 4 orthogonal viewpoints. These views are rendered using our Gaussian avatar pipeline and serve as a controlled evaluation set where ground-truth camera parameters and rendering conditions are known.
Real Captured Data.
To assess performance under realistic conditions, we collect a real-world multi-view dataset using PointGray industrial cameras. This dataset contains 5,400 frames of freely moving mice captured from four calibrated viewpoints. More details about camera setup and calibration are provided in the supplementary materials.
Baselines.
  • TripoSR (tochilkin2024triposr): A state-of-the-art open-source Large Reconstruction Model (LRM) known for strong single-view 3D generation performance.
  • Triplane-GS (zou2024triplane): A LRM variant based on Gaussian representations.
  • LGM (tang2024lgm): A multi-view diffusion-based method that reconstructs 3D Gaussian point clouds.
  • InstantMesh (xu2024instantmesh): A mesh-based 3D reconstruction method using multi-view diffusion priors.
All methods are evaluated on the task of novel view synthesis, since obtaining ground-truth 3D geometry for mouse subjects is impractical. Additionally, we include surface geometry metrics (IoU against multi-view visual hulls) in the supplementary material. We use Track-Anything (yang2023trackanythingsegmentmeets) to remove background regions in real videos.
Dataset Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
[…]
TripoSR 13.673 0.8032 0.18255
[…]
LGM 14.460 0.8805 0.13274
[…]
TripoSR 11.518 0.8114 0.19672
[…]
Table 1: Quantitative results on both synthetic and real-world datasets. MoReMouse achieves the best performance across all metrics in both domains.
Figure 6: Qualitative comparison on real data. MoReMouse maintains consistent geometry and appearance, while baselines often show pose and shape artifacts. Occasional tail discontinuities arise from imperfect Track-Anything masks under reflections or occlusions; our model remains robust and recovers plausible tail shapes.
2 Experimental Results
Quantitative Results.
Across all three metrics for novel view synthesis, MoReMouse consistently outperforms the baseline methods, demonstrating its ability to effectively infer plausible mouse geometry and pose from a single image. We also observe that the quantitative metrics on the real-world dataset are generally lower than on the synthetic dataset, especially for PSNR. This is largely due to the camera-mouse configuration: in the real dataset, the cameras are fixed while the mouse moves, whereas MoReMouse and all baselines assume that the mouse is centered and aligned across input-output views. This mismatch introduces pixel-level alignment errors, reducing PSNR. However, the perceptual metrics SSIM and LPIPS remain high, indicating that MoReMouse still reconstructs plausible and correct mouse poses and shapes. This confirms that models trained on synthetic data can generalize to real-world mouse motion capture scenarios. Full tables of IoU metrics, expanded baseline comparisons, and runtime/memory usage are provided in the supplementary.
Qualitative Results.
To visually compare MoReMouse with baseline methods, we select three representative frames from both the synthetic and real-world sequences. As shown in Fig. 5 and Fig. 6, MoReMouse generates mice with coherent appearance and physically plausible postures, accurately capturing tail and limb movements. These results remain stable even though each frame is inferred independently, showing smooth motion continuity. In contrast, baseline methods often exhibit structural distortions in novel views and inconsistent pose transitions across frames.
3 Ablation Study
Comparison with Finetuned TripoSR
To further validate the benefits of our architectural design, we fine-tune TripoSR on our mouse dataset for 120 epochs with a learning rate of $1\times 10^{-6}$. As shown in Table 2 and Fig. 7, the fine-tuned TripoSR shows noticeable improvement over its original version but still underperforms MoReMouse across most metrics. This indicates that, beyond data adaptation, our architectural choices—such as triplane encoding and geodesic embeddings—contribute significantly to reconstruction quality. We also fine-tune Triplane-GS and report its performance in the supplementary, which reveals reconstruction artifacts particularly around limbs and the tail.
Dataset Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
[…]
TripoSR-Tuned 18.162 0.9461 0.09354
Effect of Feature Embedding
Dataset Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
[…]
w/o embedding 21.805 0.9655 0.05467
[…]
w/o embedding 18.250 0.9469 0.08767
5 Discussion
Summary
In this paper, we present MoReMouse, the first monocular dense 3D reconstruction model tailored for laboratory mice. Unlike existing single-view 3D methods designed for general objects or humans, MoReMouse addresses the unique challenges posed by small animals, including non-rigid deformation, low surface texture, and limited training data. Our approach integrates transformer-based triplane representations, a Gaussian avatar modeling pipeline, and geodesic surface feature embeddings to achieve robust and high-fidelity reconstructions from a single image. We demonstrate strong performance across both synthetic and real datasets, outperforming state-of-the-art reconstruction baselines in terms of visual quality and perceptual consistency. Ablation studies further validate the architectural contributions of our model beyond dataset-specific priors. We hope that MoReMouse can serve as a foundation for future research in computational ethology and animal modeling, promoting more precise, efficient, and scalable behavioral analysis tools for biomedical research.
Limitations.
Ethical Statement
Please refer to supplementary file for Ethical Statements.
Acknowledgements
Appendix
Figure A1: Pipeline for constructing the Dense-view Dataset and synthetic evaluation data. The first 6,000 frames of the “markerless_mouse_1” 6-view sequence are used to build the dense-view dataset for training, while the last 6,000 frames are reserved for generating the synthetic evaluation set. For each training frame, AGAM is rendered from 64 randomly sampled viewpoints on a sphere of radius $2.22$, resulting in 12,00012{,}000 multi-view scenes.
Appendix A1 Ethical Statement
Appendix A2 Dataset Construction
1 Dense-view Dataset Generation
As illustrated in Fig. A1, we construct our Dense-view Dataset using the first 6,000 frames of the “markerless_mouse_1” video sequence. For each frame, we sample two independent sets of $64$ viewpoints, producing 12,00012{,}000 multi-view scenes in total, each containing $64$ images.
For viewpoint sampling, we first align the local coordinate origin of the generated Animatable Gaussian Avatar of Mouse (AGAM)—given by the origin of its internally driven mesh model—to the world coordinate origin. The Gaussian mouse model is uniformly scaled down by a factor of $180$ so that the mouse fits inside a unit sphere. We then randomly sample camera centers on a sphere of radius $2.22$, allowing for both top-down and bottom-up observations, so our dataset can accommodate to scenarios where mice are recorded from above or through a transparent floor. The rendered images have a resolution of $800\times 800$ pixels, and the FoV is fixed at $29.86^{\circ}$.
2 Synthetic Evaluation Dataset Generation
We use the last 6,000 frames from the “markerless_mouse_1” sequence to construct a Synthetic Evaluation Dataset consisting of four orthogonal viewpoints (front, back, left, and right). Following a similar placement strategy as in the Dense-view Dataset, we align the local coordinate origin of the Gaussian mouse model with the world coordinate origin and uniformly scale it down by a factor of $180$ so that it fits within a unit sphere. The four cameras are positioned on a sphere of radius $2.22$ with optical axes pointing toward the scene origin, evenly spaced in azimuth by $90^{\circ}$. The rendered images use a fixed FoV of $29.86^{\circ}$ and have a resolution of $800\times 800$.
To verify that the poses in the evaluation set differ from those in the training set, we visualize the pose distributions of the Dense-view Dataset and Synthetic Evaluation Dataset using t-SNE (maaten2008visualizing), as shown in Fig. A2. Note that the evaluation set contains a large number of unseen poses, ensuring that the reported performance reflects the model’s ability to generalize to novel poses.
Figure A2: t-SNE visualization of mouse poses. The visualization shows that the training and testing poses are separated in the embedding space. Blue points are poses from Dense-view training dataset, red points are poses from the evaluation set.
3 Real Captured Data
As shown in Fig. A3, we collect a real-world multi-view dataset using PointGrey industrial cameras. The capture setup consists of four synchronized cameras positioned around a freely moving C57BL/6 mouse. The mouse is placed inside a transparent cylinder with a diameter of 20cm and a height of 20cm. Each camera records at a resolution of $2048\times 1500$ pixels, with a horizontal field-of-view (FoV) of $22.34^{\circ}$ and a vertical FoV of $16.44^{\circ}$. This real-world dataset serves to evaluate the generalization of our model from synthetic training data to realistic environments. Before feeding the images to the network, all mice are cropped out with background removed.
Figure A3: Real-world dataset acquisition setup. Four PointGrey industrial cameras are arranged to capture a freely moving C57BL/6 mice. This dataset enables the evaluation of cross-domain generalization from synthetic to real-world conditions.
Figure A4: Qualitative comparison on the original 6-view “markerless_mouse_1” dataset. Our method produces more geometrically accurate and visually consistent reconstructions compared to other baselines. We also observe that TRELLIS occasionally outputs nearly blank images, indicating instability when handling mouse poses.
Figure A5: Qualitative comparison on the real-world multi-view dataset. Our method reconstructs anatomically coherent geometry and consistent appearance across views, whereas competing methods often produce structural distortions or inconsistent textures.
Appendix A3 More Experimental Results
In this section, we present additional experimental results beyond those reported in the main paper. In addition to the four baseline methods compared in the main text (TripoSR (tochilkin2024triposr), Triplane-GS (zou2024triplane), LGM (tang2024lgm), and InstantMesh (xu2024instantmesh)), we further compare our method against two recent reconstruction approaches:
  • 3D Fauna (li2024learning): A template-free method for learning 3D reconstruction of animals from online monocular images.
  • TRELLIS (xiang2025structured): A structured generative model designed to capture fine-grained articulations through a hierarchical representation.
We also report a surface geometry metric based on the IoU between the rendered reconstruction and multi-view visual-hull masks, offering a direct measure of geometric alignment.
To further examine the behavior of Triplane-GS on our task, we fine-tune it on our mouse dataset for 200 epochs using a learning rate of $1\times 10^{-6}$. However, the resulting Gaussian point clouds remain blurry and lack sharp geometric boundaries, making it difficult to capture fine structures such as limbs and tails.
We additionally adapt 3D-Fauna’s pan-category deformable model to our 6-view mouse video dataset via fine-tuning. During training, each iteration samples data with a probability of $0.6$ from the original 3D-Fauna dataset and $0.4$ from our mouse dataset. We freeze the base shape bank while fine-tuning the instance deformation network, pose predictor, and texture field. Optimization is performed for 500k iterations with a batch size of 10, using Adam (learning rate $1\times 10^{-4}$ for the instance network, $1\times 10^{-3}$ for the base network, weight decay 0). However, we observe that 3D-Fauna struggles to fit our mouse dataset, which covers the entire viewing sphere, leading to suboptimal performance in multi-view reconstruction.
Finally, we benchmark inference efficiency: MoReMouse runs at approximately 0.91 seconds per frame on an NVIDIA A100 GPU, consuming 7.9 GB of GPU memory during inference.
1 Comparison on Original Images of “markerless_mouse_1”
In addition to the evaluations on the Synthetic Evaluation Data constructed from the last 6,000 frames of the “markerless_mouse_1” sequence, we further evaluate all methods by directly reconstructing the same 6,000 frames from the original 6-view captures of “markerless_mouse_1”. Qualitative comparisons are shown in Fig. A4, and quantitative results are summarized in Table A1.
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ IoU $\uparrow$
[…]
TripoSR 15.837 0.9202 0.1453 0.4474
[…]
Triplane-GS 15.703 0.9182 0.1398 0.4682
Triplane-GS-Tuned 17.549 0.9293 0.1117 0.5966
LGM 14.079 0.9051 0.1519 0.3230
InstantMesh 14.876 0.9127 0.1313 0.4080
Fauna 14.261 0.9086 0.1488 0.4030
Fauna-Tuned 14.054 0.9089 0.1481 0.2956
TRELLIS 13.199 0.9146 0.1304 0.1242
Table A1: Quantitative results on the original 6-view “markerless_mouse_1” dataset. Cyan shading indicates top-3 performance.
2 Comparison on Real Data
we evaluate all compared approaches on our captured real multi-view dataset. Qualitative comparisons are presented in Figure A5, and quantitative results are summarized in Table A2. Our method consistently achieves the best performance across all three metrics, indicating superior reconstruction fidelity and perceptual quality.
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ IoU $\uparrow$
[…]
TripoSR 11.518 0.8114 0.19672 0.4741
[…]
Triplane-GS 16.789 0.9298 0.11002 0.5743
Triplane-GS-Tuned 18.628 0.9469 0.09239 0.6099
LGM 15.215 0.9197 0.12838 0.4138
InstantMesh 15.631 0.9175 0.11339 0.5466
3D Fauna 13.284 0.8837 0.12890 0.4467
3D Fauna-Tuned 13.726 0.8977 0.12038 0.4254
TRELLIS 15.087 0.9279 0.10968 0.3095
Table A2: Quantitative results on the real-world multi-view dataset. Cyan shading indicates top-3 performance.
Appendix A4 Failure cases
Figure A6 showcases typical failure cases on real-world data. The first row highlights the model’s inability to generalize to rare upright postures, due to the absence of such poses in the training set. This leads to implausible geometry and emphasizes the need for more pose-diverse avatar data.
Figure A6: Representative failure cases. Top: failure due to lack of standing mouse data in training, leading to unrealistic reconstructions. Bottom: failure due to tail-induced occlusion from shaved fur, which interferes with head feature extraction.
Appendix A5 More Details about MoReMouse Training
Our model is trained in two stages: the NeRF-based volumetric rendering stage is trained for $60$ epochs to ensure proper density field convergence, followed by the DMTet-based explicit geometry stage trained for $100$ epochs to refine surface details. The overall optimization objective combines geometric, photometric, and perceptual terms, and is defined as:
@@ 2 Comparison on Real Data
We summarize the architectural and training hyperparameters of the MoReMouse model in Table A3, covering key components such as the image encoder, triplane representation modules, multi-head MLP, NeRF and DMTet renderers, and training configurations.
@@
Hidden layers for density head 11
Hidden layers for feature/deformation heads 11
@@
Density bias 1.0-1.0
Renderer-DMTet SDF bias 4.0-4.0
@@
$\lambda_{\text{mse}}$ $1.0$
$\lambda_{\text{LPIPS}}$ $1.0$
@@
$\lambda_{\text{smooth}}$ $0.2$
$\lambda_{\text{depth}}$ $0.2$
Table A3: MoReMouse Architecture and Training Hyperparameters